Will is w years old. Ben is three years older and Jan is twice as old as Ben.
A. Write an equation, in terms of w, for Ben's age and Jan's age.
B. If you add together the ages of Will, Ben and Jan the total comes to 41 years . Write an algebraic equation and find out how old Will, Ben , and Jan are.

C. In how many years will Jan be twice as old as Will ? Explain your answer.
Please Help Needed ASAP! !! Thank you, please

Answer by Finavon(81)

Will w
Ben w+3
Jan 2(w+3)

B) w+w+3+2(w+3) = 41
Combine
4w + 9 = 41
4w = 32
w=8
Will 8
Ben 11
Jan 22

C)
n years in the future Will's age 8+n; Jan's age 22+n
(8+n) * 2 = 22 + n
16 + 2n = 22 + n
Subtract n + 16 from each side
n = 6
In 6 years time when Will is 14 and Jan is 28